5. Suppose your group would like to do a research about the weight loss on not having dinner, and you believe that not having dinner will significantly contribute on weight loss. Your group gathers a sample of data by recording the weights for each person for 16 days, and there are 25 participants. During the 16 days, the average total loss for the 25 participants are 5 pounds, with the standard deviation is 2.4. Could we conclude that weight loss is significantly contributed by no having dinner?

Please answer the following questions for 5

(hints: for questions 5, if there is a significant contribution, the number should be different from 0. Also, please be careful with the number of observations, think about if n is 16 or 25)

a. Please indicate the null hypotheses

b. Please indicate the alternative hypotheses

c. Please indicate this is a one tail test or two tail test

d. Please indicate if z table or t table will be used

e. If z table will be used, type NA for this question. If t table will be used, indicate the degree of freedom.

f. Please indicate r69orchMC7dSFISUxCDDxqjrHZFfNP41WCP++sb6, aVKdpVaZwAAAABJRU5ErkJggg==, 3t85ffzsjYcv3n3++Z+Vi19MRkXbyERXjheWPrAZand x7e2HD6oPP2BXM3NzNpDlwqiPeRHy2IcuN1grEhh

g. Please calculate the statistical value
Answer #1

The hypothesis being tested is:

H0: µ = 0

Ha: µ ≠ 0

This is a two-tailed test.

The t-table will be used.

The degree of freedom is = 25 - 1 = 24

phpJk76oP.png = 5

µ0 = 0

s = 2.4

n = 25

The test statistic, t = (x - µ)/s/√n

t = 5-0 - = 10.42

Please give me a thumbs-up if this helps you out. Thank you!